# Depth-Weighted Detection of Behaviours of Risk in People with Dementia using Cameras

### v2

@@ Abstract
The behavioural and psychological symptoms of dementia, such as agitation and aggression, present a significant health and safety risk in residential care settings. Many care facilities have video cameras in place for digital monitoring of public spaces, which can be leveraged to develop an automated behaviours of risk detection system that can alert the staff to enable timely intervention and prevent the situation from escalating. However, one of the challenges in our previous study was the presence of false alarms due to disparate importance of events based on distance. To address this issue, we proposed a novel depth-weighted loss to enforce equivalent importance to the events happening both near and far from the cameras; thus, helping to reduce false alarms. We further propose to utilize the training outliers to determine the anomaly threshold. The data from nine dementia participants across three cameras in a specialized dementia unit were used for training. The proposed approach obtained the best area under receiver operating characteristic curve performance of $0.852$, $0.81$ and $0.768$, respectively, for the three cameras. Ablation analysis was conducted for the individual components of the proposed approach and effect of frame size and frame rate. The performance of the proposed approach was investigated for cross-camera, participant-specific and sex-specific behaviours of risk detection. The proposed approach performed reasonably well in reducing false alarms. This motivates further research to make the system more suitable for deployment in care facilities.
@@ 1 Introduction
Behaviours of risk detection is a challenging problem due to their diversity and rare occurrence, making supervised classification approaches inappropriate. An anomaly detection approach is more plausible, where a method can be trained to learn the normal behaviour characteristics and identify high variation cases as anomalies during testing [3]. In our previous work [4], a spatio-temporal autoencoder was used to detect behaviours of risk as anomalies. However, there were significant issues with false positives (or false alarms), that could cause usability issues. In general, video anomaly detection (VAD) autoencoder approaches can become sensitive to the background, leading to high reconstruction errors when a major part of the background is blocked [5]. The objects or people closer to the camera cover a larger portion of the background than those further away, leading to significantly higher reconstruction errors and causing increased false positives. To tackle this problem, we propose a novel loss where the depth of the pixels is used as weight to diminish the overshadowing effect caused by the larger area of the people/objects close to camera. The aim is to reduce false positives by enforcing equivalent importance in analyzing the events happening in the unit irrespective of the distance to camera. In the absence of anomalies in the training set, we propose to utilize unusual activities that are not behaviours of risk, such as large objects or crowded scenes as proxy outliers to determine a threshold for detecting behaviours of risk. Expanding on our previous work [4] that analyzed behaviours of risk in one dementia participant using a single camera, here we extend the analysis to nine participants and three cameras. The performance for detecting behaviours of risk can vary across individuals depending upon the type of behaviours of risk, number of episodes recorded and demographic variables such as sex. To address this variability, we also conduct participant-specific and sex-specific analyses of the effectiveness of detecting behaviours of risk by the proposed approach. The main contributions of this work are:
@@
    We analyze the performance of the proposed approach for generalization across cameras, participant-specific and sex-specific behaviours of risk detection in PwD.
@@ 2 Related Work
Behaviours of Risk Detection: The current approaches for detecting behaviours of risk in PwD use different types of modalities, including wearable devices [6], computer vision [7, 4], radio waves [8] and multimodal / ambient sensing [9]. As this work uses videos, further review only discusses the studies that either use videos alone or in combination with other sensors. Qiu et al. [9] proposed a multimodal information fusion model using various sensors, including pressure sensors, ultrasound sensors, infrared sensors, video cameras, and acoustic sensors. A layered classification architecture, consisting of a hierarchical hidden Markov model and a support vector machine, was employed. However, their results were based on mock-up data generated through simulation. Chikhaoui et al. [10] developed an ensemble learning classifier to detect agitation using data from a Kinect camera and an accelerometer. Ten participants were instructed to perform six different agitated and aggressive behaviours. However, the study did not specify whether the participants were healthy volunteers or PwD. Fook et al. [7] introduced a method employing a multi-layer architecture of a probabilistic classifier based on hidden markov model and a support vector machine classifier. However, the video data involved a person in bed, and it was unclear whether the participants were healthy individuals or PwD. For all the above studies, it appears that these systems were conceptual and were never tested with real patients, so their performance is unknown in a real-world setting. Acknowledging the importance of employing real-world data, in our previous work, we proposed an unsupervised convolutional autoencoder to detect agitation in PwD using videos collected from a specialized dementia unit from a single participant and a single camera [11]. In our later work [4] with the same limited dataset, we presented privacy-protecting VAD approaches to detect behaviours of risk in PwD using body pose information for the participants or semantic segmentation masks.
Video Anomaly Detection: VAD is a task in computer vision, aimed at identifying unusual or anomalous events in video sequences in the absence of anomalous samples during training. Among these, autoencoders and depth-based methods have emerged as prominent approaches. Liu et al. [12] proposed an appearance-motion united autoencoder framework for VAD which learned the appearance and motion normality jointly. Le and Kim [13] proposed an attention-based residual autoencoder which encoded both spatial and temporal information in a unified way. Schneider et al. [14] used foreground-aware loss and investigated how existing algorithms could leverage depth modality from Kinect camera to enhance detection performance. Denkovski et al. [15] proposed a multi-objective reconstruction and prediction-based loss to detect falls using multiple camera modalities including depth within a window of sequential frames.
@@
Reducing False Positives: A VAD method can land into usability issues in the presence of increased false positives. Therefore, it is important to assess the rate of false positives (or false alarms) in these applications. We briefly review existing methods that handle false positives for VAD scenarios. Doshi and Yilmaz [16] proposed an online VAD method and a procedure for selecting an operating threshold that satisfied a desired false alarm rate. The framework was composed of deep learning-based feature extraction from video frames and a statistical sequential anomaly detection algorithm. Zhou et al. [17] proposed an anomaly trajectory detection framework for online traffic incident alerts on freeways. It used an adversarial loss to enable the autoencoder to learn a better normal trajectory pattern that is beneficial for reducing false alerts, while an abnormal trajectory discriminator was established and trained to detect small mean shifts and filter out instantaneous false alerts. Mozaffari et al. [18] proposed an online and multivariate anomaly detection method that was suitable for the timely and accurate detection of anomalies. They analyzed an asymptotic false alarm rate and provided a procedure for selecting a threshold to satisfy a desired false alarm rate. Singh et al. [19] aimed to address the challenge of a high false alarm rate resulting from the trade-off between learning to reconstruct and distinguish anomalies. They proposed a constrained generative adversarial network where white gaussian noise was added to the input to boost the robustness of the model.
@@
The existing methods tried to reduce false positives by leveraging algorithmic designs or choosing the threshold depending on a desired false positive rate. However, as described in Section 1, the sensitivity to the distance of objects/people from the camera can be a determinant factor in the increased false positives in VAD tasks, especially for scenes with large variations in depth, such as corridors or hallways. Hence, we utilize the depth of pixels to handle this issue and enforce equivalent importance to events happening far and close to the camera, thus reducing the false positives.
@@ 3.1 Data Collection
The video data used in this work was collected as part of a larger multimodal sensor study [6, 20] at the Specialized Dementia Unit (SDU) at Toronto Rehabilitation Institute, University Health Network (UHN), Toronto, Canada, between November $2017$ and October $2019$. The participants were admitted to the SDU from long-term care homes for the evaluation and management of behavioural and psychological symptoms in advanced dementia. Inclusion criteria encompassed individuals aged over 55 years, diagnosed with dementia and with a history of agitation or aggressive behaviours. The participants who expressed dissent to use of a wearable device (one of the study sensors, not part of this video-based study) were excluded from the study. Data collection ceased under either of the following circumstances: absence of documented agitation for one week, repeated removal of the wearable device, sickness, death or discharge from the unit, or upon completion of two months of data collection. The data collection was approved by the research ethics board (UHN REB#14-8483). Informed consent was obtained from substitute decision-makers for all participants. The staff also provided written consent for video recording within the unit. Additionally, all staff and decision-makers of participants shown in the images in the paper have granted written consent for the publication of their images.
@@
3.2 Selected Dataset
For this study, we selected a subset of videos, approximately $93.5$ hours for each camera, based on 30 minutes before and after each documented behaviour of risk event from the larger video dataset comprising the top three cameras (see Fig. 1). Table 1 presents the demographic information of the participants. The selected video data captured the behaviours of risk events for nine dementia participants with an age range of 6693669366–9366 – 93 years and a mean age of $81.22\pm 8.12$ years. There were three male and six female participants.
3.3 Data Annotation
Two research assistants were hired on a part-time basis and independently reviewed selected videos containing behaviour of risk events across Cam1, 2 and 3. We had the timestamps of each behaviour of risk event, so the annotators did not have to watch the entire video feed but only look at the selected segment for each behaviour of risk event. This activity took approximately two months and they annotated all the events corresponding to motor or physical agitation. Due to the lack of audio during data collection, it was not possible to annotate verbal aggression. Anvil [21] tool was used for the annotation of camera videos. Both the research assistants annotated the videos separately for the presence of behaviours of risk events and other activities, namely, the presence of unusual activities, crowded scenes and large objects. The two assistants individually annotated $305$ and $383$ behaviours of risk events across three cameras and nine participants. For the final labelling, an event was labelled as behaviour of risk if either of the research assistants annotated it as a behaviour of risk. For cases where both the assistants annotated an event as behaviour of risk but disagreed in the start/end times of the annotation, the duration with common agreement was labelled as a behaviour of risk. An inter-rater agreement analysis was performed, where Cohen’s kappa [22], Krippendorff alpha [23] and percentage agreement metrics were calculated over the annotations of both research assistants. The metrics were calculated based on the timestamps (in seconds) of the annotations of the two annotators. This was done to measure the agreeability of the two annotations over the categorization of the events into normal and behaviours of risk events. The Cohen’s kappa, Krippendorff alpha and percentage agreement were determined as $0.48$, $0.48$ and $0.99$, respectively. The value of Cohen’s kappa and Krippendorff alpha can range from $-1$ to $1$, where $1$ and $-1$ represent a perfect agreement and disagreement, and 00 represents agreement expected from random chance.
3.4 Data Preprocessing
The original resolution, frame rate and bit depth of the videos were $352\times 240$, $30$ frames per second (fps) and $24$, respectively. The frames were sampled from the videos at $15$ fps, selecting every second frame. The frames were then converted to grayscale, normalized to range $[0,1]$ (pixel values divided by $255$) and resized to $64\times 64$ resolution using the python programming language. This was done to minimize the computational expenses associated with trainable parameters. The frames were then stacked separately to form non-overlapping $5$-second windows ($75$ frames per window) (based on previous work [11]). The labels for behaviours of risk were refined in accordance with the $5$-second windows. Training and test sets were generated individually for the three cameras. The training sets were composed of normal behaviours from daily activities only. The test sets were composed of both normal and behaviours of risk events. The training set size was fixed around an average of $1224.41$ minutes across three cameras. The test set consists of a combination of behaviours of risk events from nine participants across three cameras along with normal behaviours. In the test sets, the available duration of annotated behaviours of risk events was limited. Hence, the duration of normal behaviour in test sets was chosen to keep a similar imbalance ratio (as in previous work [4]), which is approximately $7\%$ for behaviours of risk events. Table 2 shows the size of training and test sets for each of the three cameras. The test set was used to evaluate the performance of methods for detecting behaviours of risk. The number of $5$-second windows and types of behaviours of risk across participants and cameras in the test set are presented in Table 3.
@@ 3.5 Behaviours of Risk Detection
In our previous work [11, 4], we used a customized spatio-temporal convolutional autoencoder (3DCAE) to detect behaviours of risk from data from one participant and a single camera. 3DCAE learned to reconstruct the input videos with normal behaviour during training. Hence, during testing, normal instances were expected to have a lower reconstruction error, while instances with high reconstruction error were treated as anomalies (behaviours of risk in our case). However, it showed high false positives due to sensitivity to activities such as crowded scenes, large objects, or other events happening near the camera due to reasons described in Section 1. Such instances are extremely common in normal day-to-day activities happening in care units. As such, there was a need to handle this issue to make the method suitable for real-world deployment. Hence in this work, we propose a novel depth-weighted loss, where the depth of the pixels is used as a weight factor to enforce equivalent importance to both behaviours of risk and non-behaviours of risk events happening far and near the camera. It is hypothesized that the weighted-depth loss will lead to a reduction in false positives in this problem. We also propose a new way to determine the threshold for detecting behaviours of risk as anomalies using the abnormal non-behaviours of risk activities in the training set as outliers. Hence, the proposed approach consists of the novel depth-weighted loss and utilizing the training outliers to determine the threshold.
@@
[2pt]Refer to caption(b)
Figure 2: (a) Illustration of the geometric relationship between a 3D prism (side length $L$) and the area of its projection in the image plane. (b) Two identical 3D prisms placed at different distances from camera position $C$ ($Z_{a}>Z_{b}$) produce a smaller projected area ($A_{a}<A_{b}$).
@@
where $f$ is the focal length and $Z$ is the distance from the $3$D object to the camera position. Assuming the difference in depth $Z$ between these vertices is negligible ($Z\approx Z_{1}\approx Z_{2}$), the projections on the image plane according to equation (1) are:
@@
Consider an anomaly event occurring at different distances $Z_{a}$ and $Z_{b}$ from the camera. As it is the same event with same volume, it will have same value of $\overline{AS}_{\text{pixel}}$ and length $L$ at both distances. As illustrated in Fig. 2 (b), when projected in the image plane, this results in different area values $A_{a}$ and $A_{b}$ in the image plane, with $A_{a}<A_{b}$ if $Z_{a}>Z_{b}$. Since both anomalies have the same value of $\overline{AS}_{\text{pixel}}$, according to equation (4), the difference in area will result in a different anomaly score. To ensure depth invariance, we need the anomaly to have the same score $AS_{inv}$ regardless of its distance from the camera. This is achieved by introducing a factor $K$:
$AS_{inv}=K_{a}\cdot AS_{A_{a}}=K_{b}\cdot AS_{A_{b}}$ (5)
Combining the equations (3), (4) and (5), we can express $AS_{inv}$ as :
$AS_{inv}=K_{a}\overline{AS}_{\text{pixel}}\left(fL\right)^{2}\frac{1}{Z_{a}^{2}}=K_{b}\overline{AS}_{\text{pixel}}\left(fL\right)^{2}\frac{1}{Z_{b}^{2}}$ (6)
@@
Therefore, given the constant value of $\overline{AS}_{\text{pixel}}$ and $\left(fL\right)^{2}$, to achieve depth invariance, the factors $K_{a}$ and $K_{b}$ must satisfy the following conditions:
$K_{a}=Z_{a}^{2},\quad K_{b}=Z_{b}^{2}$ (7)
@@
Based on this, we propose the following novel loss where depth per pixel is used as weight during the calculation of reconstruction error as mean squared error to enable the methods to enforce equivalent importance to both near and far pixels.
$L_{dmse}=\frac{1}{N_{e}}\sum_{l=1}^{W}\sum_{i=1}^{S}\sum_{j=1}^{S}{\color[rgb]{0,0,0}\phi_{l,i,j}}\>\>(I_{l,i,j}-O_{l,i,j})^{2}$ (8)
@@
where, $I$ is the input pixel, $O$ is the reconstructed pixel, $S$ is the spatial size, and $W$ represents the number of frames in a window. $N_{e}=W\times S\times S$ is the total number of pixels in a window. The depth-based weight hyperparameter $\phi$ can be expressed as a function of $Z^{2}$, i.e., $\phi=\mathfrak{f}(Z^{2})$. The value of $\phi$ is chosen using cross-validation from among different functions of $Z^{2}$, namely $\left(Z^{2},Z^{2}/2,\sqrt{Z^{2}}=Z\right)$, by leveraging the proxy validation set in a similar way as the threshold. The weighted-depth loss helped the method to treat both the near and far behaviours of risk events with similar importance, which in turn helped to reduce false positives. The dense prediction transformer [25] was used to estimate the depth map during both training and testing. During testing, the depth-weighted loss was used as anomaly score for detecting behaviours of risk.
@@
Threshold Determination: In an anomaly detection approach, it can be challenging to determine an operating threshold due to the absence of anomalous samples in the training set. As discussed in Section 3.3, the data was annotated for other unusual activities, such as large objects, crowded scenes and other infrequent activities in the training set. Due to the unusual nature of these annotated activities relative to the activities of daily living in a dementia unit, we propose to treat these activities as proxy outliers to create a proxy validation set to determine the threshold. Following a similar approach as in our previous work [26], we chose each reconstruction error as a threshold during training and found the corresponding F1-score on the proxy validation set. Finally, the reconstruction error with the highest F1-score was chosen as the operating threshold. The operating threshold was then utilized during testing to detect behaviours of risk in the test set and the results were reported in Section 4.
@@
Experimental Setup: Transformers and variational autoencoders have performed well in the existing literature for VAD [27, 28]. In this work, we have used three different anomaly detection methods, namely, 3DCAE, a self-supervised masked convolutional transformer block-based autoencoder (CTBAE) and a variational autoencoder (VAE) for evaluation. The 3DCAE method has been adapted from our previous work [4]. The CTBAE method used a self-supervised masked convolutional transformer block [27] embedded after the encoder layer and its output was passed as input to the decoder. VAE [29] learned to generate two vectors that represent the parameters (mean and variance) of a distribution from which the latent vector was sampled. For fairness during comparison, the encoder and decoder architecture of 3DCAE, CTBAE and VAE were kept the same and they used interquartile range (IQR) generated proxy-validation set [26] for threshold calculation. The proposed depth-weighted approach was integrated as part of the 3DCAE, CTBAE and VAE methods and their depth variants are referred to as depCAE, depCTBAE and depVAE, respectively. The depCAE method followed an encoder-decoder architecture, where the encoder consisted of $2$D convolution and max-pooling layers, with a kernel size of ($1\times 3\times 3$), stride ($1\times 1\times 1$) and padding ($0\times 1\times 1$), followed by batch normalization and ReLU operations. The decoder used $2$D deconvolution layers ($2$D transposed convolution operation) with a kernel size of ($1\times 3\times 3$), strides ($1\times 1\times 1$), ($1\times 2\times 2$), ($1\times 2\times 2$) and paddings ($0\times 1\times 1$), ($0\times 1\times 1$), ($0\times 1\times 1$) for first, second, and third $2$D deconvolution layers, respectively. The overall architecture of the depCAE is presented in Fig. 3. In the depCAE method, $W$ = 75 and $N_{e}$ = $75\times 64\times 64=307200$.
Evaluation Metrics: The proposed approach aims to reduce false positives; however, it is also important to consider false negatives when detecting behaviours of risk in PwD. Hence, when evaluating performance, it is important to consider a metric that can give a balanced estimate of the performance on both false positive rate (FPR) and false negative rate (FNR). Precision and recall can be expressed as,
$Precision=\frac{TPR}{TPR+\frac{1}{\gamma}FPR}$ (9)
$Recall=\frac{TPR}{TPR+FNR}$ (10)
where TPR refers to true positive rate and $\gamma=\frac{No.\>of\>positives\>(P)}{No.\>of\>negatives\>(N)}$ is class ratio. As seen in eqns. 9 and 10, precision and recall give an estimate of the performance over FPR and FNR, respectively. F1-score is the harmonic mean of precision and recall, and provides a reasonable performance estimate over TPR, FPR and FNR, simultaneously. Matthew’s Correlation Coefficient (MCC) gives an overall performance estimate over true positives (TP), true negatives (TN), false positives (FP) and false negatives (FN) and is expressed as,
$MCC=\frac{TP*TN-FP*FN}{\sqrt{(TP+FP)*P*N*(TN+FN)}}$ (11)

### v3

@@ Abstract
Background: The behavioural and psychological symptoms of dementia, such as agitation and aggression, present a significant health and safety risk in residential care settings. Many care facilities have video cameras in place for digital monitoring of public spaces, which can be leveraged to develop an automated behaviours of risk detection system that can alert the staff to enable timely intervention. However, one of the challenges is the presence of false alarms in these systems. Methods: We proposed a novel depth-weighted loss to enforce equivalent importance to the events happening both near and far from the cameras, aimed to reduce false alarms. We further propose to utilize the training outliers to determine the anomaly threshold. The data from nine dementia participants across three cameras in a specialized dementia unit were used for training. Ablation analysis was conducted for the individual components of the proposed approach and the effect of frame size and frame rate. The performance of the proposed approach was investigated for cross-camera, and participant-specific behaviours of risk detection. Results: The proposed approach obtained the best area under receiver operating characteristic curve performance of $0.852$, $0.81$ and $0.768$, respectively, for the three cameras. The proposed approach recorded a lower false positive rate than the existing methods, indicating its efficacy in reducing false alarms. Significance: This study motivates further research to make the behaviours of risk detection system more suitable for deployment in care facilities, improving the health and safety of people with dementia.
@@ 1 Introduction
Behaviours of risk detection is a challenging problem due to their diversity and rare occurrence, making supervised classification approaches inappropriate. An anomaly detection approach is more plausible, where a method can be trained to learn the normal behaviour characteristics and identify high variation cases as anomalies during testing [3]. In our previous work [4], a spatio-temporal autoencoder was used to detect behaviours of risk as anomalies. However, there were significant issues with false positives (or false alarms), that could cause usability issues. In general, video anomaly detection (VAD) autoencoder approaches can become sensitive to the background, leading to high reconstruction errors when a major part of the background is blocked [5]. The objects or people closer to the camera cover a larger portion of the background than those further away, leading to significantly higher reconstruction errors and causing increased false positives. To tackle this problem, we propose a novel loss where the depth of the pixels is used as weight to diminish the overshadowing effect caused by the larger area of the people/objects close to camera. The aim is to reduce false positives by enforcing equivalent importance in analyzing the events happening in view irrespective of the distance to camera. In the absence of anomalies in the training set, we propose to utilize unusual activities that are not behaviours of risk, such as large objects or crowded scenes as proxy outliers to determine a threshold for detecting behaviours of risk. Expanding on our previous work [4] that analyzed behaviours of risk in one dementia participant using a single camera, here we extend the analysis to nine participants and three cameras. The performance for detecting behaviours of risk can vary across individuals depending upon the type of behaviours of risk, and number of episodes recorded. To address this variability, we also conduct participant-specific analysis of the effectiveness of detecting behaviours of risk by the proposed approach. The main contributions of this work are:
@@
    We analyze the performance of the proposed approach for generalization across cameras, and participant-specific behaviours of risk detection in PwD.
@@ 2 Related Work
Behaviours of Risk Detection: The current approaches for detecting behaviours of risk in PwD use different types of modalities, including wearable devices [6], computer vision [7, 4], radio waves [8] and multimodal / ambient sensing [9]. As this work uses videos, further review only discusses the studies that either use videos alone or in combination with other sensors. Qiu et al. [9] proposed a multimodal information fusion model using various sensors, including pressure sensors, ultrasound sensors, infrared sensors, video cameras, and acoustic sensors. A layered classification architecture, consisting of a hierarchical hidden Markov model and a support vector machine, was employed. However, their results were based on mock-up data generated through simulation. Chikhaoui et al. [10] developed an ensemble learning classifier to detect agitation using data from a Kinect camera and an accelerometer. However, the study did not specify whether the participants were healthy volunteers or PwD. Fook et al. [7] introduced a method employing a multi-layer architecture of a probabilistic classifier based on hidden markov model and a support vector machine classifier. However, the video data involved a person in bed, and it was unclear whether the participants were healthy individuals or PwD. For all the above studies, it appears that these systems were conceptual and were never tested with real patients, so their performance is unknown in a real-world setting. Acknowledging the importance of employing real-world data, we proposed an unsupervised convolutional autoencoder to detect agitation in PwD using videos collected from a specialized dementia unit from a single participant and a single camera [11]. In our later work [4] with the same limited dataset, we presented privacy-protecting VAD approaches using body poses or semantic segmentation masks to detect behaviours of risk in PwD.
@@
Reducing False Positives: A VAD method can land into usability issues in the presence of increased false positives. Therefore, it is important to assess the rate of false positives (or false alarms) in these applications. We briefly review existing methods that handle false positives for VAD scenarios. Doshi and Yilmaz [12] proposed an online VAD method and a procedure for selecting an operating threshold that satisfied a desired false alarm rate. The framework was composed of deep learning-based feature extraction from video frames and a statistical sequential anomaly detection algorithm. Zhou et al. [13] proposed an anomaly trajectory detection framework for online traffic incident alerts on freeways. It used an adversarial loss to enable the autoencoder to learn a better normal trajectory pattern that is beneficial for reducing false alerts. Mozaffari et al. [14] proposed an online and multivariate anomaly detection method that was suitable for the timely and accurate detection of anomalies. They analyzed an asymptotic false alarm rate and provided a procedure for selecting a threshold to satisfy a desired false alarm rate. Singh et al. [15] proposed a constrained generative adversarial network where white gaussian noise was added to the input to boost the robustness of the model.
@@
The existing methods tried to reduce false positives by leveraging algorithmic designs or choosing the threshold depending on a desired false positive rate. However, as described in Section 1, the sensitivity to the distance of objects/people from the camera can be a determining factor in the increased false positives in VAD tasks, especially for scenes with large variations in depth, such as corridors or hallways. Hence, we utilize the depth of pixels to handle this issue and enforce equivalent importance to events happening far and close to the camera, thus reducing the false positives.
@@ 3.1 Data Collection
The video data used in this work was collected as part of a larger multimodal sensor study [6, 16] at the Specialized Dementia Unit (SDU) at Toronto Rehabilitation Institute, University Health Network (UHN), Toronto, Canada, between November $2017$ and October $2019$. Inclusion criteria encompassed individuals aged over 55 years, diagnosed with dementia and with a history of agitation or aggressive behaviours. Data collection ceased under either of the following circumstances: absence of documented agitation for one week, sickness, death or discharge from the unit, or upon completion of two months of data collection. The data collection was approved by the research ethics board (UHN REB#14-8483). Informed consent was obtained from substitute decision-makers for all participants. The staff also provided written consent for video recording within the unit. For this study, we selected a subset of videos, approximately $93.5$ hours for each of the three cameras (see Fig. 1), based on 30 minutes before and after each documented behaviour of risk event. Table 3 presents the demographic information of the participants. The selected video data captured the behaviours of risk events for nine dementia participants with an age range of $66$-$93$ years and a mean age of $81.22\pm 8.12$ years. There were three male and six female participants.
@@
3.2 Data Annotation
Two research assistants independently reviewed selected videos containing behaviour of risk events across Cam 1, 2 and 3, and annotated all the events corresponding to motor or physical agitation. Due to the lack of audio during data collection, it was not possible to annotate verbal aggression. Anvil [17] tool was used for the annotation of camera videos. Both the research assistants also annotated the videos for other activities, namely, the presence of unusual activities, crowded scenes and large objects. The two assistants individually annotated $305$ and $383$ behaviours of risk events across three cameras and nine participants. For the final labelling, an event was labelled as behaviour of risk if either of the research assistants annotated it as a behaviour of risk. For cases where both the assistants annotated an event as behaviour of risk but disagreed in the start/end times of the annotation, the duration with common agreement was labelled as a behaviour of risk. An inter-rater agreement analysis was performed, where Cohen’s kappa [18], Krippendorff alpha [19] and percentage agreement metrics were calculated over the timestamps of the annotations of both research assistants. The Cohen’s kappa, Krippendorff alpha and percentage agreement were determined as $0.48$, $0.48$ and $0.99$, respectively. The value of Cohen’s kappa and Krippendorff alpha can range from $-1$ to $1$, where $1$ and $-1$ represent a perfect agreement and disagreement, and 00 represents agreement expected from random chance.
3.3 Data Preprocessing
The original resolution, frame rate and bit depth of the videos were $352\times 240$, $30$ frames per second (fps) and $24$, respectively. The frames were sampled from the videos at $15$ fps, converted to grayscale, normalized to range $[0,1]$ (pixel values divided by $255$) and resized to $64\times 64$ resolution. This was done to minimize the computational expenses associated with trainable parameters. The frames were then stacked separately to form non-overlapping $5$-second windows ($75$ frames per window) (based on previous work [11]). Training and test sets were generated individually for the three cameras. The training sets were composed of normal behaviours from daily activities only. The test set consisted of a combination of behaviours of risk events from nine participants and normal behaviours. Table 3 shows the size of training and test sets for each of the three cameras. The test set was used to evaluate the performance of methods for detecting behaviours of risk. The number of $5$-second windows and types of behaviours of risk across participants and cameras in the test set are presented in Table 3.
@@ 3.4 Behaviours of Risk Detection
In our previous work [11, 4], we used a customized spatio-temporal convolutional autoencoder (3DCAE) to detect behaviours of risk from data from one participant and a single camera. 3DCAE learned to reconstruct the input videos with normal behaviour during training. Hence, during testing, normal instances were expected to have a lower reconstruction error, while instances with high reconstruction error were treated as anomalies (behaviours of risk in our case). However, it showed high false positives due to sensitivity to activities such as crowded scenes, large objects, or other events happening near the camera due to reasons described in Section 1. As such, there was a need to handle this issue to make the method suitable for real-world deployment. Hence in this work, we propose a novel depth-weighted loss, where the depth of the pixels is used as a weight factor to enforce equivalent importance to both behaviours of risk and non-behaviours of risk events happening far and near the camera. It is hypothesized that the weighted-depth loss will lead to a reduction in false positives. The proposed approach consists of the novel depth-weighted loss and utilizing the training outliers to determine the threshold.
@@
Figure 2: (a) Illustration of the geometric relationship between a 3D prism (side length $L$) and the area of its projection in the image plane. (b) Two identical 3D prisms placed at different distances from camera position $C$ produce different projected areas. In this example, if $Z_{a}>Z_{b}$, then $A_{a}<A_{b}$.
@@
where $f$ is the focal length and $Z$ is the distance from the $3$D object to the camera position. Assuming the difference in depth $Z$ between these vertices is negligible ($Z\approx Z_{1}\approx Z_{2}$), the projections on the image plane according to eq. (1) are:
@@
Consider an anomaly event occurring at different distances $Z_{a}$ and $Z_{b}$ from the camera. As it is the same event with same volume, it will have same value of $\overline{AS}_{\text{pixel}}$ and length $L$ at both distances. As illustrated in Fig. 2 (b), when projected in the image plane, this results in different area values $A_{a}$ and $A_{b}$ in the image plane, with $A_{a}<A_{b}$ if $Z_{a}>Z_{b}$. Since both anomalies have the same value of $\overline{AS}_{\text{pixel}}$, according to eq. (4), the difference in area will result in a different anomaly score. To ensure depth invariance, we need the anomaly to have the same score regardless of its distance from the camera. This is achieved by introducing a factor $K$:
$K_{a}\cdot AS_{A_{a}}=K_{b}\cdot AS_{A_{b}}$ (5)
Performing square root on both sides and combining the equations (3), (4) and (5), we can express eq. (5) as :
$\sqrt{K_{a}\overline{AS}_{\text{pixel}}\left(fL\right)^{2}\frac{1}{Z_{a}^{2}}}=\sqrt{K_{b}\overline{AS}_{\text{pixel}}\left(fL\right)^{2}\frac{1}{Z_{b}^{2}}}$
$\implies K_{a}^{{}^{\prime}}\sqrt{\overline{AS}_{\text{pixel}}}fL\frac{1}{Z_{a}}=K_{b}^{{}^{\prime}}\sqrt{\overline{AS}_{\text{pixel}}}fL\frac{1}{Z_{b}}$ (6)
@@
Therefore, given the constant value of $\sqrt{\overline{AS}_{\text{pixel}}}$ and $fL$, to achieve depth invariance, the factors $K_{a}^{{}^{\prime}}$ and $K_{b}^{{}^{\prime}}$ must satisfy the following conditions:
$K_{a}^{{}^{\prime}}=Z_{a},\quad K_{b}^{{}^{\prime}}=Z_{b}$ (7)
@@
Based on eq. 7, we propose the following novel loss where depth $Z$ is used as weight during the calculation of reconstruction error as mean squared error to enable the methods to enforce equivalent importance to both near and far pixels.
$L_{dmse}=\frac{1}{N_{e}}\sum_{l=1}^{W}\sum_{i=1}^{S}\sum_{j=1}^{S}Z_{l,i,j}\>\>(I_{l,i,j}-O_{l,i,j})^{2}$ (8)
@@
where, $Z$ is the depth weight, $I$ is the input pixel, $O$ is the reconstructed pixel, $S$ is the spatial size, and $W$ represents the number of frames in a window. $N_{e}=W\times S\times S$ is the total number of pixels in a window. The weighted-depth loss helped the method to treat both the near and far behaviours of risk events with similar importance, which in turn helped to reduce false positives. The dense prediction transformer [21] was used to estimate the depth map during both training and testing. During testing, the depth-weighted loss was used as anomaly score for detecting behaviours of risk.
@@
Threshold Determination: In an anomaly detection approach, it can be challenging to determine an operating threshold due to the absence of anomalous samples in the training set. As discussed in Section 3.2, the data was annotated for other unusual activities, such as large objects, crowded scenes and other infrequent activities in the training set. We propose to treat these activities as proxy outliers to create a proxy validation set to determine the threshold. We first create two sets: inliers ($I$) and proxy outliers ($O$). The set $I$ is composed of all the training windows. The set $O$ is comprised of windows annotated as other activities. Thereafter, $I$ and $O$ are further divided into (90%-10% split): training ($I_{t}$, $O_{t}$) and validation ($I_{v}$, $O_{v}$) sets [22] (see Figure 4). A method is trained on $I_{t}$, and the reconstruction error of all the windows in $I_{t}$ is determined. Then, each reconstruction error is chosen as an intermediate threshold, and the corresponding F1-score is calculated using the proxy validation set, which is the combination of $I_{v}$ and $O_{v}$. The reconstruction error with the highest F1-score on the proxy validation set is chosen as the threshold for testing.
@@
Experimental Setup: Transformers and variational autoencoders have performed well in the existing literature for VAD [23, 24]. In this work, we have used three different anomaly detection methods, namely, 3DCAE, a self-supervised masked convolutional transformer block-based autoencoder (CTBAE) and a variational autoencoder (VAE) for evaluation. The 3DCAE method has been adapted from our previous work [4]. The CTBAE method used a self-supervised masked convolutional transformer block [23] embedded after the encoder layer and its output was passed as input to the decoder. VAE [25] learned to generate two vectors that represent the parameters (mean and variance) of a distribution from which the latent vector was sampled. For fairness during comparison, the encoder and decoder architecture of 3DCAE, CTBAE and VAE were kept the same and they used interquartile range (IQR) generated proxy-validation set [22] for threshold calculation. The proposed depth-weighted approach was applied as part of the 3DCAE, CTBAE and VAE methods and their depth variants are referred to as depCAE, depCTBAE and depVAE, respectively. The depCAE method followed an encoder-decoder architecture, where the encoder consisted of $2$D convolution and max-pooling layers, with a kernel size of ($1\times 3\times 3$), stride ($1\times 1\times 1$) and padding ($0\times 1\times 1$), followed by batch normalization and ReLU operations. The decoder used $2$D deconvolution layers ($2$D transposed convolution operation) with a kernel size of ($1\times 3\times 3$), strides ($1\times 1\times 1$), ($1\times 2\times 2$), ($1\times 2\times 2$) and paddings ($0\times 1\times 1$), ($0\times 1\times 1$), ($0\times 1\times 1$) for first, second, and third $2$D deconvolution layers, respectively. The overall architecture of the depCAE is presented in Fig. 4.